# Point Linguist Model: Segment Any Object via Bridged Large 3D-Language Model

### v1

@@ Abstract
Index Terms:
Point Cloud; Large Language Model; Object Segmentation
@@ I Introduction
Recent advancements in Multi-modal Large Language Models (MLLMs) for point cloud understanding [1, 2, 3, 4, 5, 6, 7, 8] have introduced a new interaction paradigm, enabling natural language-based communication with point cloud scenes. These pioneers have inspired efforts for the 3D object segmentation tasks [9, 10, 11, 12, 13, 14] with LLMs. Despite the success in 2D MLLMs [15, 16, 17, 18, 19], without a segmentation priori like SAM [20], successful paradigms like LISA [15] cannot be replicated.
@@
To address these issues, we propose Point Linguist Model (PLM) that bridges the representation gap between LLMs and dense 3D point clouds with our Object-centric Discriminative Representation (OcDR) and Geometric Reactivation Decoder (GRD). Firstly, our OcDR use the object-centric (OC) tokens as the visual input of the LLM. As seen in the left panel of Figure 1, the OC token naturally maintains object-wise distinctions, allowing the LLM to directly access object-oriented features. The straightforward information injection helps the LLM to better identify different objects and their relationships in the scene. Secondly, based on the captured object relationships, we propose a distractor-supervised mechanism to refine object differentiation. As shown in Figure 2, it selects hard negative distractors—objects with semantic proximity to the target—and incorporates them as additional supervision signals. By explicitly supervising the target with these distractors during training, our model enhances object identity discrimination, leading to more accurate and robust segmentation. For accurate segmentation, our GRD preserves comprehensive dense features throughout the pipeline. As shown in the right panel of Figure 1 (b), we merge the comprehensive dense features into the OC tokens and preserve scene details during the LLM’s reasoning pipeline. In the decoding process, the LLM-inferred geometry with corresponding dense features is reactivated through an attention mechanism before propagating to the final output mask.
@@ II-B Multi-modal Large Language Models
Multi-modal Large Language Models (MLLMs) have garnered significant attention from researchers for their powerful ability to transfer the advanced reasoning capabilities of LLMs into the visual domain. Foundation models in both 2D [37, 38, 39, 40] and 3D [5, 6, 41, 8] have focused on aligning image/point cloud and text features through instruction tuning, laying the foundation for vision Large Language Models. Recent approaches have increasingly focused on the more challenging task of dense, pixel-level prediction, e.g., LISA [15] advances the 2D object segmentation by integrating the Segmentation Anything Model (SAM) [20] and LLaVA [37]. Such a paradigm shows that MLLMs can derive segmentation capacity from a priori model like SAM. In the 3D domain, SegPoint [22] also built a unified framework for various segmentation tasks, combining pre-trained Uni3D [42] with the LLaMA model [43]. However, lacking a foundational segmentation model like SAM in the 3D domain, dense prediction requires MLLMs to learn scene parsing implicitly—a process that demands massive paired data [44, 45] and is therefore data-inefficient and suboptimal.
@@ III-B2 Multi-Modal Large Language Model
During LLM processing, the [point] token will be padded by the $\mathbf{f}_{OC}$. As illustrated in Figure 3, after recognizing the [SEG] token, the LLM will output a predicted token for segmentation within the output sequence $\mathbf{y}_{pred}$, donated as $\mathbf{y}_{\texttt{[SEG]}}$. The predicted $\mathbf{y}_{\texttt{[SEG]}}$ is then fed into the next stage for mask generation. We adopt the LLaMA2-7B [43] as the LLM and apply LoRA [46] for efficient fine-tuning on the LLM.
@@ III-B3 Geometric Reactivation Decoder
The decoder processes mask queries $\mathbf{y}_{\texttt{Mask}}$ through a sequence of attention layers to generate output queries $\hat{\mathbf{y}}_{\texttt{Mask}}$, which are used for object masks. The mask queries first go through a multi-head self-attention mechanism, allowing interaction among queries to refine their representations. The queries are then passed through two multi-head cross-attention layers. The first cross-attention layer attends to the LLM output $\mathbf{y}_{\texttt{[SEG]}}$ (as keys and values), integrating compact cross-modal features and retrieving the object referenced by the language instruction. The second cross-attention layer attends to the OcDR $\mathbf{f}_{OC}$. By revisiting the $\mathbf{f}_{OC}$, the decoder reactivates the preserved details information of the whole scene, identifying those surrounding the target object and reinforcing inter-object relationships. The decoder is constructed by stacking these layers $n$ times. This process can be formalized as follows:
@@ III-C Supervision with Distractors & Matching Loss
Specifically, the generated output masks from our GRD will then be matched with the ground truth of both the target objects and our designed distractors. The distractors are defined as objects with semantic proximity to the target, i.e., those objects with the same or akin semantic category of the targets. For instance, when there is more than one Chair in the scene, or similar categories, like Bed and Sofa Bed are present in one scene. We extract these distractors from the semantic categories of ScanNet200 [10] and construct a semantic distractor dataset based on open-vocabulary and referring segmentation benchmarks.
@@ III-D Training Pipeline
Our model is trained end-to-end with prompts that supervise both open-vocabulary segmentation and referring segmentation, ensuring robust generalization across diverse object categories and language instructions:
TABLE I: Prompt template for the language model.
User: [point] Can you segment the bookshelf in this point cloud? or User: [point] Where is the object: a bookshelf in the corner that is next to the door. Can you segment the described object? Assistant: Sure, segmented [SEG].
@@
We aggregate essential training data from the referring point cloud segmentation benchmark, including ScanRefer [12], ReferIt3D [13] (both Nr3D and Sr3D), and Multi3DRefer [14]. We also aggregate ScanNet200 [10] from the open vocabulary segmentation benchmark, which also serves as the foundational scene dataset for the aforementioned referring benchmarks. Following previous works on 3D segmentation [22, 36], we adopt mIoU as the primary metric, defined as the average of all per-point cloud scene Intersection-over-Union (IoU). Additionally, we report AP@25 and AP@50, which are based on IoU thresholds of 25% and 50%, respectively, measuring the accuracy of instance predictions at different levels of overlap with ground truth.
@@ IV Experiments
TABLE II: Open Vocabulary Instance Segmentation results on SacnNetV2 [9] and S3DIS [11]. B/N represent Base and Novel class splitting in the open vocabulary setting. We follow the standard split presented in previous methods [52, 53, 27].
@@
TABLE III: Open vocabulary semantic segmentation result in ScanNetV2 [9] and ScanNet200 [10]. We report the mIoU result.
@@
TABLE IV: Open vocabulary semantic segmentation on different novel class settings. We report the mIoU result of the standard split following previous methods.
@@
IV-A Compare to State-of-the-Art
Open-Vocabulary Instance Segmentation
We compare our method with existing 3D open-vocabulary models in the same setting. It is worth noting that the performance on S3DIS is achieved in a zero-shot setting, without scenes seen during training. For a fair comparison, we conduct OVIS while excluding the “other furniture” class from ScanNetV2 and the “clutter” class from S3DIS, as their definitions are ambiguous. As a result, ScanNetV2 includes 17 classes, and S3DIS contains 12 classes. We follow the category splits from previous methods [52, 36] and compare performance on novel classes. As shown in Table II, our method demonstrates impressive performance in OVIS, surpassing all previous methods in all 3 novel class settings. This proves the effectiveness of language instruction with the LLMs. Specifically, PLM consistently outperforms the prior state-of-the-art method, OpenIns3D [36], across all benchmarks.
On ScanNetV2, PLM achieves the highest AP50 and AP25 scores in all settings. Even in the most challenging 17-class setting, PLM shows a strong lead with 38.4% AP50 and 46.2% AP25, outperforming OpenIns3D by $\uparrow$9.7% and $\uparrow$7.3%, respectively. In a zero-shot way, similar trends are observed in the S3DIS dataset, where PLM consistently achieves the highest performance in all settings. In the 12-class setting, PLM secures the top position with 29.3% AP50 and 35.4% AP25, showing a 1.0% and 5.9% improvement over OpenIns3D.
We also provide the per-class results in Section IV-C
@@ IV-A Compare to State-of-the-Art
To further validate the effectiveness of our proposed PLM, we evaluate its performance on the OVSS task, following the standard evaluation protocols from prior works [27, 28, 25]. We compare our PLM with the latest state-of-the-art method Diff2Scene [25] and XMask3D [27]. Table III presents the mIoU results on ScanNetV2 [9] and ScanNet200 [10] datasets. Compared to previous methods, PLM demonstrates superior performance across both datasets, achieving 66.0% mIoU on ScanNetV2 and 43.5% mIoU on ScanNet200, significantly surpassing previous state-of-the-art approaches. Table IV further breaks down the OVSS results across various novel class splits. We observe that PLM consistently achieves the highest performance across all settings, demonstrating its strong generalization ability in recognizing novel classes in open-vocabulary settings.
Referring Expression Segmentation
Table V compares the results on RES and GRES. Previous LLM-based methods like SegPoint [22] only output a single segmentation mask. It manages the GRES by combining all target masks as one, resulting in suboptimal performance due to the lack of instance-level differentiation. In contrast, our PLM enables flexible segmentation of multiple objects, achieving state-of-the-art results (42.1% mIoU) on MultiRefer3D and significantly surpassing SegPoint by $\uparrow$6.0% and other previous methods, demonstrating its superior ability to understand multi-object references. As for the single-object Referring Expression Segmentation (RES), PLM also demonstrates strong performance across benchmarks. It achieves 43.1% mIoU on ScanRefer, slightly below the previous expert model RefMask3D (44.8%), while still maintaining the lead on Nr3D (32.4%) and Sr3D (34.9%), surpassing all previous methods.
TABLE V: Referring segmentation result. We conduct the generalized referring expression segmentation on the MultiRefer3D [14] dataset. And conduct the single-object referring expression segmentation on the ScanRefer [12] and ReferIt3D [14] datasets.
@@
As shown in Figure 4 (i), PLM is also capable of reasoning and comprehending implicit user instruction, just like LISA [15]. However, our study is constrained by the lack of a standardized benchmark for reasoning expression segmentation. While some works [22, 62, 63] have introduced related datasets, they remain closed-source as of this writing. Therefore, we qualitatively demonstrate the reasoning capabilities of PLM, with additional samples provided in the Figure 5.
@@ IV-B Ablation Study
After extensively demonstrating our superiority across various tasks, we conduct comprehensive ablation studies to validate the rationality of the OcDR and our design choices. We select the challenging MultiRefer3D and ScanRefer benchmarks as our ablation baselines.
OcDR with distractor-supervised
We compare our OcDR with existing methods that rely on dense point patches, which necessitate heavy pre-alignment processing. We utilize the pretrained Uni3D [42], a backbone widely adopted in prior work [57, 45].
Comparing the performance on the downstream tasks, results in Table VI reveal that using Uni3D directly leads to a significant performance drop (-6.5% on MultiRefer3D, -5.3% on ScanRefer), confirming that our structured OcDR provides better object grounding and segmentation quality. While increasing the token length from 512 to 1024 slightly improves ScanRefer performance (+0.9%), the results remain far below our setting, highlighting the limitations of relying solely on excessive tokenization without structured OcDR.
To show the superiority of the design of our OcDR, we also compare it with another alternative setting: Isolated object tokens generated by applying Uni3D on the mask proposal from Mask3D, a paradigm widely used in previous OC-based methods [45, 44]. Similarly, the second setting (mask proposal + Uni3D, in Table VI) also suffers from degraded performance. While leveraging Uni3D on mask proposals provides a structured object-centric representation, it lacks holistic scene awareness, leading to a notable performance drop.
TABLE VI: Ablation on OcDR with different settings.
@@
TABLE VII: Per-class results of OVIS on ScanNetV2. Performance on novel classes is highlighted. The AP50 result is reported.
@@
TABLE VIII: Per-class results of OVIS on S3DIS. Performance on novel classes is highlighted. The AP50 result is reported.
@@
TABLE IX: Comparison of data efficiency with 50% of training data.
Methods Multi3DRefer ScanRefer
SegPoint [22] 36.1 41.7
PLM (50% data) 40.9 41.2
TABLE X: Comparison of computation cost on different representations.
Representations inferencing time (s) acceptable batch size
OcDR (150 tokens) 0.28 12
Uni3D [42] (512 tokens) 0.51 12
Uni3D [42] (1024 tokens) 1.07 4
TABLE XI: Ablation study on different parameter settings for the GRD. Blue line represents the default settings: Decoder nn=3; mask queries = 16.
GRD Multi3DRefer ScanRefer
default settings 42.1 43.1
w/o Decoder 27.9 32.5
Decoder nn=1 40.5 39.8
Decoder nn=5 42.1 42.8
mask queries = 4 41.1 40.7
mask queries = 24 41.7 42.8
mask queries = 48 41.0 38.6
w/o bbox loss 40.8 42.4
Our OcDR also demonstrates superior data efficiency. We train our PLM using just 50% of the training data and benchmark it against Segpoint [22], the current SOTA point-patch token–based LLM model. Results are shown in Table IX. We match Segpoint’s performance using only half the training data, and even surpass it on the Multi3DRefer [14] multi-object segmentation task. Furthermore, we evaluate the model’s efficiency, comparing the OcDR with the point-patch tokens. We use excessive point-patch tokens generated by Uni3D [42], a strong point cloud encoder pre-trained with language alignment. Uni3D provides a rich semantic understanding of 3D objects and enables better alignment with LLMs. It serves as the point cloud encoder in the SegPoint [22] model. As shown in Table X, Uni3D incurs a significant computational cost, leading to slower inference and reduced batch efficiency. Increasing token length from 512 to 1024 further exacerbates the issue, more than doubling the inference time while drastically reducing the acceptable batch size. These two results highlight the inefficiency of relying on excessive point-patch tokens, making it impractical for scalable deployment.
GRD with distractor-supervised
As shown in Table XI, our decoder plays a crucial role in segmentation accuracy, significantly enhancing multi-object segmentation and instance separation.
Finally, we tested the depth of the decoder, the length of mask queries, and the bounding box loss in our GRD. We select the optimal parameters based on consistently robust results. Results shown in Table XI. We select the optimal parameters based on consistently robust results.

### v2

@@ I Introduction
Recent advancements in Multi-modal Large Language Models (MLLMs) for point cloud understanding [60, 50, 64, 43, 62, 17, 52, 44] have introduced a new interaction paradigm, enabling natural language-based communication with point cloud scenes. These pioneers have inspired efforts for the 3D object segmentation tasks [7, 48, 3, 5, 1, 67] with LLMs. Despite the success in 2D MLLMs [31, 55, 46, 61, 66, 14, 13, 34, 9], without a segmentation priori like SAM [28], successful paradigms like LISA [31] cannot be replicated.
@@
To address these issues, we propose Point Linguist Model (PLM) that bridges the representation gap between LLMs and dense 3D point clouds with our Object-centric Discriminative Representation (OcDR) and Geometric Reactivation Decoder (GRD). Firstly, our OcDR uses the object-centric (OC) tokens as the visual input of the LLM. As seen in the left panel of Figure 1, the OC token naturally maintains object-wise distinctions, allowing the LLM to directly access object-oriented features. The straightforward information injection helps the LLM to better identify different objects and their relationships in the scene. Secondly, based on the captured object relationships, we propose a distractor-supervised mechanism to refine object differentiation. As shown in Figure 2, it selects hard negative distractors—objects with semantic proximity to the target—and incorporates them as additional supervision signals. By explicitly supervising the target with these distractors during training, our model enhances object identity discrimination, leading to more accurate and robust segmentation. For accurate segmentation, our GRD preserves comprehensive dense features throughout the pipeline. As shown in the right panel of Figure 1 (b), we merge the comprehensive dense features into the OC tokens and preserve scene details during the LLM’s reasoning pipeline. In the decoding process, the LLM-inferred geometry with corresponding dense features is reactivated through an attention mechanism before propagating to the final output mask.
@@ II-B Multi-modal Large Language Models
Multi-modal Large Language Models (MLLMs) have garnered significant attention from researchers for their powerful ability to transfer the advanced reasoning capabilities of LLMs into the visual domain. Foundation models in both 2D [37, 69, 2, 53] and 3D [62, 17, 36, 44] have focused on aligning image/point cloud and text features through instruction tuning, laying the foundation for vision Large Language Models. Recent approaches have increasingly focused on the more challenging task of dense, pixel-level prediction, e.g., LISA [31] advances the 2D object segmentation by integrating the Segmentation Anything Model (SAM) [28] and LLaVA [37]. Such a paradigm shows that MLLMs can derive segmentation capacity from an a priori model like SAM. In the 3D domain, SegPoint [15] also built a unified framework for various segmentation tasks, combining pre-trained Uni3D [68] with the LLaMA model [54]. However, lacking a foundational segmentation model like SAM in the 3D domain, dense prediction requires MLLMs to learn scene parsing implicitly—a process that demands massive paired data [38, 19] and is therefore data-inefficient and suboptimal.
@@ III-B2 Multi-Modal Large Language Model
During LLM processing, the [point] token will be padded by the $\mathbf{f}_{OC}$. As illustrated in Figure 3, after recognizing the [SEG] token, the LLM will output a predicted token for segmentation within the output sequence $\mathbf{y}_{pred}$, donated as $\mathbf{y}_{\texttt{[SEG]}}$. The predicted $\mathbf{y}_{\texttt{[SEG]}}$ is then fed into the next stage for mask generation. We adopt the LLaMA2-7B [54] as the LLM and apply LoRA [18] for efficient fine-tuning.
@@ III-B3 Geometric Reactivation Decoder
The decoder processes mask queries $\mathbf{y}_{\texttt{Mask}}$ through a sequence of attention layers to generate output queries $\hat{\mathbf{y}}_{\texttt{Mask}}$, which are used for object masks. The mask queries first go through a multi-head self-attention mechanism, allowing interaction among queries to refine their representations. The queries are then passed through two multi-head cross-attention layers. The first cross-attention layer attends to the LLM output $\mathbf{y}_{\texttt{[SEG]}}$ (as keys and values), integrating compact cross-modal features and retrieving the object referenced by the language instruction. The second cross-attention layer attends to the OcDR $\mathbf{f}_{OC}$. By revisiting the $\mathbf{f}_{OC}$, the decoder reactivates the preserved details information of the whole scene, identifying those surrounding the target object and reinforcing inter-object relationships. The decoder is constructed by stacking these layers $n$ times, which can be formalized as follows:
@@ III-C Supervision with Distractors & Matching Loss
Specifically, the generated output masks from our GRD will then be matched with the ground truth of both the target objects and our designed distractors. The distractors are defined as objects with semantic proximity to the target, i.e., those objects with the same or akin semantic category as the targets. For instance, when there is more than one Chair in the scene, or similar categories, like Bed and Sofa Bed are present in one scene. We extract these distractors from the semantic categories of ScanNet200 [48] and construct a semantic distractor dataset based on open-vocabulary and referring segmentation benchmarks.
@@
Additional details of the loss definition are provided in the APPENDIX. A
@@ III-D Training Pipeline
Our model is trained end-to-end with prompts that supervise both open-vocabulary segmentation and referring segmentation, ensuring robust generalization across diverse object categories and language instructions. We design sets of task-specific prompt templates, shown in Table I. During training, one prompt is randomly sampled from the corresponding prompt set at each iteration to increase linguistic diversity. During evaluation, we use the first one in each list for all methods to ensure stable results.
TABLE I: Prompt template for PLM training.
Semantic Segmentation: [point] Can you segment the {class_name} category in this point cloud? [point] Please segment the {class_name} category in this point cloud. [point] What is {class_name} category in this point cloud? Please respond with segmentation mask. [point] What is {class_name} category in this point cloud? Please output segmentation mask. Referring Segmentation: [point] With a description: {description} Please respond with segmentation mask. [point] Giving the referring sentence: {description} Please output segmentation mask. [point] Where is the object: {description} Can you segment the described object? [point] In this scene: {description} Please segment the described object.
@@
We aggregate essential training data from the referring point cloud segmentation benchmark, including ScanRefer [5], ReferIt3D [1] (both Nr3D and Sr3D), and Multi3DRefer [67]. We also aggregate ScanNet200 [48] from the open vocabulary segmentation benchmark, which also serves as the foundational scene dataset for the aforementioned referring benchmarks. Following previous works on 3D segmentation [15, 22], we adopt mIoU as the primary metric, defined as the average of all per-point cloud scene Intersection-over-Union (IoU). Additionally, we report AP@25 and AP@50, which are based on IoU thresholds of 25% and 50%, respectively, measuring the accuracy of instance predictions at different levels of overlap with ground truth. Our baselines were trained and evaluated under identical data augmentations and prompt settings to ensure a fair comparison. Specifically, we adopt a unified augmentation pipeline for all training, including random point dropout, rotating and flipping, scaling, jittering, and color augmentations. Noticed that we did not apply any test-time augmentation in our experiments.
@@
Distractor design
For the distractor setting, we use two straightforward designs to generate the distractor list for each sample. Firstly, for the semantic segmentation task, distractors are constructed by identifying semantically similar categories within the predefined class set (e.g., Bed vs. Sofa Bed, Office chair vs. Armchair). When objects from such semantically overlapping categories occur in the same scene, they are treated as distractors. Noted that this process is mutual. We provide further examples in Table II. Secondly, for the referring segmentation task, distractors are constructed by identifying the object with the same category as the targets. We provide an example in Figure 2, among the presented 10 chairs, only one is the target, and the other are distractors.
TABLE II: Examples for semantic distractors.
Target categorie Distractor categories
shower wall closet wall
office chair armchair, sofa chair, folded chair
tissue box mailbox
@@ IV Experiments
IV-A Compare to State-of-the-Art
Open-Vocabulary Instance Segmentation
We compare our method with existing 3D open-vocabulary models in the same setting. It is worth noting that the performance on S3DIS is achieved in a zero-shot setting, without scenes seen during training. For a fair comparison, we conduct OVIS while excluding the “other furniture” class from ScanNetV2 and the “clutter” class from S3DIS, as their definitions are ambiguous. As a result, ScanNetV2 includes 17 classes, and S3DIS contains 12 classes. We follow the category splits from previous methods [10, 22] and compare performance on novel classes. As shown in Table III, our method demonstrates impressive performance in OVIS, surpassing all previous methods in all 3 novel class settings. This proves the effectiveness of language instruction with the LLMs. Specifically, PLM consistently outperforms the prior state-of-the-art method, OpenIns3D [22], across all benchmarks.
TABLE III: Open Vocabulary Instance Segmentation results on SacnNetV2 [7] and S3DIS [3]. B/N represent Base and Novel class splitting in the open vocabulary setting. We follow the standard split presented in previous methods [10, 65, 56].
@@ IV-A Compare to State-of-the-Art
TABLE IV: Open vocabulary semantic segmentation result in ScanNetV2 [7] and ScanNet200 [48]. We report the mIoU result.
@@
TABLE V: Open vocabulary semantic segmentation on different novel class settings. We report the mIoU result of the standard split following previous methods.
@@
On ScanNetV2, PLM achieves the highest AP50 and AP25 scores in all settings. Even in the most challenging 17-class setting, PLM shows a strong lead with 38.4% AP50 and 46.2% AP25, outperforming OpenIns3D by $\uparrow$9.7% and $\uparrow$7.3%, respectively. In a zero-shot way, similar trends are observed in the S3DIS dataset, where PLM consistently achieves the highest performance in all settings. In the 12-class setting, PLM secures the top position with 29.3% AP50 and 35.4% AP25, showing a 1.0% and 5.9% improvement over OpenIns3D. We also provide the per-class results in Section IV-C
@@
To further validate the effectiveness of our proposed PLM, we evaluate its performance on the OVSS task, following the standard evaluation protocols from prior works [56, 42, 71]. We compare our PLM with the latest state-of-the-art method Diff2Scene [71] and XMask3D [56]. Table IV presents the mIoU results on ScanNetV2 [7] and ScanNet200 [48] datasets. Compared to previous methods, PLM demonstrates superior performance across both datasets, achieving 66.0% mIoU on ScanNetV2 and 43.5% mIoU on ScanNet200, significantly surpassing previous state-of-the-art approaches. Table V further breaks down the OVSS results across various novel class splits. We observe that PLM consistently achieves the highest performance across all settings, demonstrating its strong generalization ability in recognizing novel classes in open-vocabulary settings.
TABLE VI: Referring segmentation result. We conduct the generalized referring expression segmentation on the MultiRefer3D [67] dataset. And conduct the single-object referring expression segmentation on the ScanRefer [5] and ReferIt3D [67] datasets.
@@
Referring Expression Segmentation
Table VI compares the results on RES and GRES. Previous LLM-based methods like SegPoint [15] only output a single segmentation mask. It manages the GRES by combining all target masks as one, resulting in suboptimal performance due to the lack of instance-level differentiation. In contrast, our PLM enables flexible segmentation of multiple objects, achieving state-of-the-art results (42.1% mIoU) on MultiRefer3D and significantly surpassing SegPoint by $\uparrow$6.0% and other previous methods, demonstrating its superior ability to understand multi-object references. As for the single-object Referring Expression Segmentation (RES), PLM also demonstrates strong performance across benchmarks. It achieves 43.1% mIoU on ScanRefer, slightly below the previous expert model RefMask3D (44.8%), while still maintaining the lead on Nr3D (32.4%) and Sr3D (34.9%), surpassing all previous methods.
@@
As shown in Figure 4 (i), PLM is also capable of reasoning and comprehending implicit user instruction, just like LISA [31]. However, our study is constrained by the lack of a standardized benchmark for reasoning expression segmentation. While some works [15, 20, 27] have introduced related datasets, they remain closed-source as of this writing. Therefore, we qualitatively demonstrate the reasoning capabilities of PLM, with additional samples provided in the APPENDIX. B.
@@ IV-B Ablation Study
After extensively demonstrating our superiority across various tasks, we conduct comprehensive ablation studies to validate the rationality of our design choices for the OcDR and GRD. We select the challenging MultiRefer3D and ScanRefer benchmarks as our ablation baselines.
Ablation on OcDR tokens
Our OcDR encodes the entire scene in an object-centric manner at the feature level, emphasizing high-level semantic structure rather than precise mask boundaries. The subsequent GRD module refines segmentation masks only at the final stage. As a result, even when some proposals are inaccurate or missing, the overall scene representation remains informative, mitigating sensitivity to proposal errors produced by Mask3D. We analyze the robustness of PLM to the number of OcDR tokens. As shown in Table VII, although the encoder is pretrained with 150 tokens, varying the token number to 50 or 200 results in small performance changes. Specifically, using fewer tokens may omit some valid object proposals and lead to a less complete scene representation, while using more tokens may introduce noisier proposals. This indicates that OcDR does not rely on a narrowly tuned token budget and that moderate mismatch between pretraining and inference token counts does not critically affect representation quality. The model maintains stable performance, demonstrating robustness to variations in proposal quality.
TABLE VII: Compare OcDR with other point cloud encoding methods. Blue line represents the default settings of our PLM.
@@
TABLE VIII: Comparison of data efficiency with 50% of training data.
Methods Multi3DRefer ScanRefer
SegPoint [15] 36.1 41.7
PLM (50% data) 40.9 41.2
Ablation on OcDR Design
We compare our OcDR with two existing point cloud encoding methods used in the SegPoint [32] and Chat-Scene [19].
Firstly, SegPoint [32] relies on dense point patches generated by the pretrained Uni3D backbone [68]. Comparing the performance on the downstream tasks, results in Table VII reveal that using the SegPoint method with Uni3D directly leads to a significant performance drop (-6.5% on MultiRefer3D, -5.3% on ScanRefer), confirming that our structured OcDR provides better object grounding and segmentation quality. While increasing the token length from 512 to 1024 slightly improves ScanRefer performance (+0.9%), the results remain far below our setting, highlighting the limitations of relying solely on excessive tokenization without structured OcDR.
Secondly, Chat-Scene [19] uses isolated object tokens generated by applying Uni3D [68] on the mask proposal from Mask3D [49]. Similarly, the second setting (mask proposal + Uni3D, in Table VII) also suffers from degraded performance. While leveraging Uni3D on mask proposals provides a structured object-centric representation, it lacks holistic scene awareness, leading to a notable performance drop.
Ablation on Distractor Supervision
@@
TABLE IX: Comparison of computation cost on different representations. Blue line represents the default settings of our PLM.
Representations inferencing time (s) acceptable batch size
OcDR (150 tokens) 0.28 12
Uni3D [68] (512 tokens) 0.51 12
Uni3D [68] (1024 tokens) 1.07 4
TABLE X: Ablation study on different parameter settings for the GRD. Blue line represents the default settings: Decoder nn=3; mask queries = 16.
GRD Multi3DRefer ScanRefer
default settings 42.1 43.1
w/o Decoder 27.9 32.5
Decoder nn=1 40.5 39.8
Decoder nn=5 42.1 42.8
mask queries = 4 41.1 40.7
mask queries = 24 41.7 42.8
mask queries = 48 41.0 38.6
w/o bbox loss 40.8 42.4
Ablation on Data Efficiency
Our OcDR also demonstrates superior data efficiency. We train our PLM using just 50% of the training data and benchmark it against Segpoint [15], the current SOTA point-patch token–based LLM model. Results are shown in Table VIII. We match Segpoint’s performance using only half the training data, and even surpass it on the Multi3DRefer [67] multi-object segmentation task. Furthermore, we evaluate the model’s efficiency, comparing the OcDR with the point-patch tokens. We use excessive point-patch tokens generated by Uni3D [68], a strong point cloud encoder pre-trained with language alignment. Uni3D provides a rich semantic understanding of 3D objects and enables better alignment with LLMs. It serves as the point cloud encoder in the SegPoint [15] model. As shown in Table IX, Uni3D incurs a significant computational cost, leading to slower inference and reduced batch efficiency. Increasing token length from 512 to 1024 further exacerbates the issue, more than doubling the inference time while drastically reducing the acceptable batch size. These two results highlight the inefficiency of relying on excessive point-patch tokens, making it impractical for scalable deployment.
Ablation on GRD Design
As shown in Table X, our decoder plays a crucial role in segmentation accuracy, significantly enhancing multi-object segmentation and instance separation.
We also analyze the GRD design choices, including the decoder depth, mask query number, and bounding box loss, and select the default configuration based on consistently robust performance (Table X). The mask query number implicitly limits the maximum number of distractors participating in matching; when this limit is exceeded, distractors are randomly subsampled. As shown in Table X, PLM remains stable across different mask query settings, supporting the robustness of the distractor-supervised matching and the GRD design.
TABLE XI: Per-class results of OVIS on ScanNetV2. Performance on novel classes is highlighted. The AP50 result is reported.
@@
TABLE XII: Per-class results of OVIS on S3DIS. Performance on novel classes is highlighted. The AP50 result is reported.
@@
TABLE XIII: Computational performance of our PLM.
Setting Computation (TFLOPs) Throughput Memory (GB)
forward backward total samples/sec latency (ms)
Decoder nn=3 47.19 94.37 141.56 3.2 312.76 17.4
Decoder nn=5 48.79 97.57 146.36 3.1 326.79 17.4
Ablation on Hungarian Matching
We then do ablation on the Hungarian loss weights. The Hungarian matching involves five loss terms, including the classification loss, mask focal loss, Dice loss, L1 loss, and GIoU loss. We adopt the empirical weight setting from Mask3D without specific fine-tuning, and additionally evaluate a uniform-weight variant. The results are reported in Table XIV.
TABLE XIV: Ablation on Hungarian loss weight settings. The loss weights are reported in the order of classification loss, mask focal loss, Dice loss, L1 loss, and GIoU loss.
loss weights Multi3DRefer ScanRefer
2,2,5,1,2 42.1 43.1
1,1,1,1,1 39.4 40.0
As expected, the loss weights have a noticeable impact on performance, since the accuracy of Hungarian matching directly affects how well the target object is distinguished from distractors. This behavior is inherent to matching-based frameworks rather than specific to our method. Importantly, our key contribution lies not in tuning these weights, but in introducing distractor-aware matching to explicitly differentiate target objects from semantically similar distractors.